# ReflecSched: Solving Dynamic Flexible Job-Shop Scheduling via LLM-Powered Hierarchical Reflection

### v2

@@ 6.4 Token Consumption Analysis
The results in Table 4 reveal a trade-off dictated by problem scale. For Normal scale instances, which are longer and more complex, ReflecSched is significantly more token-efficient than the baseline. This is because the reflection module’s generated experience allows subsequent decision-making prompts to be much leaner. Conversely, for Small scale instances, the upfront computational cost of the initial reflection outweighs these savings, resulting in higher token consumption. This highlights an intentional architectural trade-off: ReflecSched allocates a greater budget to its initial reflection phase, an investment that proves effective for larger, more complex problems by improving solution quality while simultaneously reducing the total tokens required.
@@ Appendix B Case Study: An Analysis of the Comprehension Limitations in Smaller LLMs
This appendix provides a qualitative analysis of the reasoning process exhibited by a smaller-scale model (Qwen3-4B) when functioning as the LLM-Direct scheduler [yang2025qwen3technicalreport]. The subsequent examples, extracted directly from execution logs, are presented to illustrate specific failure modes. These examples highlight the model’s difficulties in consistently and accurately interpreting the complex, natural-language representation of the DFJSP state. These case studies serve as empirical evidence for the central claim of our main paper: that the direct application of LLMs to this scheduling problem exhibits significant limitations. This analysis, therefore, underscores the necessity for a more structured framework like ReflecSched.
@@ D.3 GEN-Bench: The Fair Evaluation Benchmark
A robust measure of the spread of the central 50% of the data [tukey1977exploratory]. Let $Q1(\mathcal{P}_{I})$ and $Q3(\mathcal{P}_{I})$ be the first and third quartiles.
@@
This quantifies the diversity of the performance ordering [jun2007normalized]. Let $p_{k}$ be the proportion of rules achieving rank $k$. The normalized Shannon entropy is:

### v3

@@ 6.7 Efficiency and Runtime Analysis
The results in Table 4 reveal a trade-off dictated by problem scale. For Normal scale instances, which are longer and more complex, ReflecSched is significantly more token-efficient than the baseline. This is because the reflection module’s generated Strategic Experience allows subsequent decision-making prompts to be much leaner. Conversely, for Small scale instances, the upfront computational cost of the initial reflection outweighs these savings, resulting in higher token consumption. However, considering the significantly improved makespan ratio achieved by ReflecSched, this highlights an intentional architectural trade-off: ReflecSched allocates a greater budget to its initial reflection phase, an investment that proves effective for larger, more complex problems by improving solution quality while simultaneously reducing the total computational and financial tokens required.
@@ Appendix B Case Study: An Analysis of the Comprehension Limitations in Smaller LLMs
This appendix provides a qualitative analysis of the reasoning process exhibited by a smaller-scale model (Qwen3-4B) when functioning as the LLM-Direct scheduler [60]. The subsequent examples, extracted directly from execution logs, are presented to illustrate specific failure modes. These examples highlight the model’s difficulties in consistently and accurately interpreting the complex, natural-language representation of the DFJSP state. These case studies serve as empirical evidence for the central claim of our main paper: that the direct application of LLMs to this scheduling problem exhibits significant limitations. This analysis, therefore, underscores the necessity for a more structured framework like ReflecSched.
@@ B.2 Misinterpretation of State Descriptors
While redesigning state descriptors to distinguish between current status and post-action predictions might seem like a viable solution to prevent such conflations, the failure observed here is structural rather than merely linguistic. The tendency of the model to misinterpret predicted metrics as historical facts reflects an inherent limitation in maintaining a consistent temporal context during one-step generation. ReflecSched addresses this challenge by providing Strategic Experience to serve as a top-down interpretive framework. This mechanism ensures that the model focuses on the strategic implications of a state, such as the potential to complete a critical job, instead of being misled by the semantic ambiguity of raw numerical descriptors.
@@ B.3 Inconsistent Adherence to Procedural Rules
The ReflecSched framework systematically mitigates the inconsistent rule adherence observed in this case study through its hierarchical architecture. While smaller models in the LLM-Direct baseline may generate contradictory rationales or overlook high-priority labels, ReflecSched implements a robust fallback and validation protocol. At each decision point, the candidate actions are filtered by the environment kernel to ensure they comply with technological constraints. If the model’s output is found to be inconsistent with the set of feasible, high-priority actions, the framework utilizes a majority-voting scheme and a heuristic-driven tie-break mechanism to restore the correct decision path. Additionally, the periodic invocation of the reflection module upon the arrival of an emergency job forces the model to re-evaluate its strategy based on the most recent evidence, thereby reinforcing the high-priority rules in the Strategic Experience prompt.
@@ D.3 GEN-Bench: The Fair Evaluation Benchmark
A robust measure of the spread of the central 50% of the data [51]. Let $Q1(\mathcal{P}_{I})$ and $Q3(\mathcal{P}_{I})$ be the first and third quartiles.
@@
This quantifies the diversity of the performance ordering [26]. Let $p_{k}$ be the proportion of rules achieving rank $k$. The normalized Shannon entropy is:
@@ E.2 The Reflection Prompt
Appendix F Data for Motivational Analysis Example
This appendix provides the complete problem instance data corresponding to the motivating example discussed in Section 4.3 and illustrated in Figure 1. Table F.1 details the candidate machines and corresponding processing times for each operation, derived directly from the schedule visualisations.
Table F.1: Processing time data for the illustrative example in Figure 1. The pairs $(M_{k},p)$ indicate that machine $k$ can process the operation in $p$ time units.
Job Operation Candidate Machines & Processing Times
Job 1 $O_{1,1}$ $(M_{2},1.90),(M_{3},1.34)$
$O_{1,2}$ $(M_{3},1.98)$
$O_{1,3}$ $(M_{2},1.98)$
Job 2 $O_{2,1}$ $(M_{1},1.59),(M_{2},2.62)$
$O_{2,2}$ $(M_{2},1.71)$
$O_{2,3}$ $(M_{1},1.94),(M_{2},1.20)$
$O_{2,4}$ $(M_{1},0.96)$
Job 3 $O_{3,1}$ $(M_{1},1.08)$
$O_{3,2}$ $(M_{3},2.22)$
Appendix G Granular Performance Records for Per-Instance Comparative Analysis
This appendix provides the complete record of raw performance results for every problem instance across the three benchmarking suites evaluated in this study, enabling a transparent comparison between ReflecSched and all baseline algorithms.
Table G.1: Comparative results of Makespan RPD (%) across different problem scales on GEN-Bench. Relative Percentage Deviation (RPD) values are expressed as percentages. Bold values indicate the best performance for each instance across all evaluated methods.
Instance Baseline Methods ReflecSched (Proposed)
GP HMPSAC IDDQN DAN PPO-OC Q3-8b Q3-14b Q3-32b DS-V3 DS-V3.2 GPT-4o GPT-5n
Small-1 74.19 11.07 12.48 32.99 34.88 20.29 10.68 0.51 13.66 0.00 2.21 0.00
Small-2 23.47 28.21 28.21 7.12 15.88 2.17 7.66 9.21 7.66 0.00 7.66 7.66
Small-3 98.11 23.97 1.96 22.54 25.95 0.00 6.39 10.60 10.60 10.60 10.60 10.60
Small-4 58.65 10.15 33.63 0.00 9.59 5.13 4.33 1.82 4.73 21.88 1.22 10.54
Small-5 25.86 17.74 19.34 35.76 35.76 6.97 0.00 19.23 10.75 9.43 9.43 9.43
Small-6 43.96 7.60 19.76 8.06 11.99 9.20 0.00 4.59 4.33 7.59 9.20 9.20
Small-7 40.63 0.00 11.75 6.52 6.52 16.50 9.36 15.43 16.73 24.95 16.49 20.42
Small-8 19.90 17.63 28.28 7.80 16.76 4.23 0.15 0.00 4.00 0.76 3.62 0.76
Small-9 29.12 21.24 6.79 6.79 2.63 1.76 1.22 3.14 1.76 0.00 2.63 6.79
Small-10 13.53 0.00 0.00 0.00 12.63 16.04 26.22 2.81 13.80 10.95 17.48 26.22
Small-11 36.07 9.18 13.53 0.00 4.15 16.09 16.06 21.05 21.34 21.05 21.05 21.05
Small-12 28.70 17.47 5.31 6.21 0.00 4.60 0.95 7.17 3.72 12.42 3.08 3.08
Small-13 46.35 0.00 13.64 9.20 12.75 16.42 3.98 7.21 7.21 7.21 8.83 7.21
Small-14 1.72 3.92 6.35 0.00 7.72 8.57 3.93 3.93 3.93 3.93 3.93 3.93
Small-15 12.07 0.00 4.94 28.44 7.28 3.87 5.94 1.29 0.00 0.00 2.58 0.00
Small-16 59.52 0.00 20.07 20.07 24.93 14.38 15.43 7.24 13.59 11.03 12.72 12.84
Small-17 32.55 14.90 14.23 22.92 35.83 2.42 0.00 3.96 0.62 16.57 3.06 3.75
Small-18 55.27 29.09 0.00 23.68 42.22 4.24 4.24 4.24 5.86 9.10 4.24 4.24
Normal-1 51.78 15.19 26.14 8.34 11.38 4.61 0.00 13.97 9.11 4.79 4.45 5.48
Normal-2 60.73 16.79 5.13 0.00 6.77 4.14 7.29 4.37 3.35 4.73 1.98 16.72
Normal-3 73.72 8.01 0.00 10.39 17.73 4.70 4.83 8.01 6.58 4.62 6.57 1.97
Normal-4 46.75 3.89 2.36 4.63 1.38 0.00 5.25 7.07 7.82 6.18 8.77 6.18
Normal-5 46.89 9.45 0.00 4.28 5.15 1.21 2.29 8.59 5.30 0.15 3.70 2.48
Normal-6 55.35 1.26 3.14 0.00 5.02 0.91 0.87 4.38 1.25 14.83 3.87 16.16
Normal-7 94.23 19.61 6.59 6.04 9.23 6.30 5.97 6.93 7.30 0.00 6.37 3.68
Normal-8 95.64 13.98 7.58 4.58 0.00 3.62 3.58 4.84 0.62 2.14 4.27 8.80
Normal-9 50.58 5.35 0.05 0.31 6.08 2.87 0.00 0.06 3.76 6.72 7.30 3.52
Normal-10 47.94 12.83 12.28 25.88 17.60 15.53 8.13 8.67 10.69 9.80 21.38 0.00
Normal-11 46.58 17.55 15.34 3.66 7.27 11.56 12.56 9.26 6.59 8.58 7.52 0.00
Normal-12 129.32 23.97 11.08 13.41 2.78 5.84 6.90 2.09 4.79 3.67 0.00 4.39
Normal-13 75.27 0.00 11.70 9.71 7.92 8.57 5.45 8.14 6.55 14.86 6.71 4.00
Normal-14 27.44 1.62 0.42 12.29 5.27 4.45 0.00 9.57 13.03 0.20 12.09 26.26
Normal-15 74.02 7.69 2.42 9.67 11.31 4.47 6.13 4.83 2.39 9.41 8.47 0.00
Normal-16 79.02 18.33 17.61 0.00 23.35 2.63 7.62 4.58 14.66 1.86 1.51 7.49
Normal-17 52.80 0.00 3.62 10.13 13.36 3.90 3.95 9.30 2.52 9.40 4.64 13.15
Normal-18 38.77 12.95 9.25 16.26 13.87 1.56 8.10 2.59 2.64 10.65 0.00 3.28
Normal-19 110.60 9.77 7.95 0.00 20.28 16.14 11.82 11.04 18.47 13.95 12.24 11.08
Normal-20 127.35 7.62 14.05 30.47 18.68 7.85 14.18 9.17 12.87 0.00 9.41 12.92
[…]
Table G.2: Makespan RPD (%) performance comparison on the established MK-Bench (Brandimarte-based). RPD values are expressed as percentages to underscore the performance variance. Bold values indicate the best performance for each instance across all evaluated methods.
Instance Baseline Methods ReflecSched (Proposed)
GP HMPSAC IDDQN DAN PPO-OC Q3-8b Q3-14b Q3-32b DS-V3.2 GPT-5n
MK01 27.76 2.84 2.63 18.68 0.00 5.05 10.68 7.32 15.95 7.39
MK02 16.93 12.67 9.40 9.46 28.15 17.32 30.32 34.32 0.00 12.67
MK03 45.19 12.02 27.47 26.44 21.15 8.17 40.38 44.71 0.00 23.56
MK04 123.64 3.34 10.16 4.02 3.76 1.74 7.11 0.70 0.00 10.16
MK05 25.71 25.79 7.63 9.09 3.35 11.79 0.00 14.14 3.47 6.25
MK06 147.97 22.42 17.22 28.05 29.73 12.49 15.66 9.33 0.00 9.33
MK07 64.92 17.44 19.30 23.03 45.40 39.18 0.00 10.90 1.28 32.31
MK08 7.72 6.10 6.25 23.46 7.29 0.00 9.74 9.43 20.13 5.96
MK09 21.80 21.49 13.68 11.51 17.77 4.88 29.11 0.00 18.70 13.68
MK10 76.36 19.06 4.71 24.35 2.49 2.36 0.00 7.46 8.80 10.61
[…]
Table G.3: Makespan RPD (%) performance comparison on JMS-Bench. RPD values are expressed as percentages to highlight subtle performance variations. Bold values indicate the best performance for each instance across all evaluated methods. All ReflecSched variants utilize the hierarchical reflection mechanism.
Instance Baseline Methods ReflecSched (Proposed)
GP HMPSAC IDDQN DAN PPO-OC Q3-8b Q3-14b Q3-32b DS-V3.2 GPT-5n
instance_01 25.58 0.00 6.47 20.51 3.67 9.30 14.88 3.67 14.88 2.55
instance_02 77.49 4.36 3.05 13.55 4.03 1.81 3.72 0.00 0.84 0.27
instance_03 30.69 0.70 0.00 3.53 7.72 11.66 5.31 9.95 18.27 10.50
instance_04 61.76 11.03 13.15 32.18 32.88 12.97 0.00 10.21 11.76 13.15
instance_05 62.60 19.04 5.01 20.33 11.79 0.00 9.35 11.79 8.88 22.90
instance_06 31.25 12.50 3.13 9.72 0.00 6.25 12.50 12.50 6.25 7.84
instance_07 7.20 0.00 1.70 18.20 15.50 9.95 2.75 3.38 3.93 0.80
instance_08 45.05 20.51 15.49 0.00 11.72 7.79 5.31 9.15 7.69 14.28
instance_09 78.09 10.69 13.38 18.73 12.90 4.96 0.00 10.63 5.29 10.36
instance_10 48.30 18.33 18.70 18.88 9.17 0.00 7.97 6.11 7.24 18.91
[…]
Appendix H Hyperparameter Settings
This appendix delineates the comprehensive hyperparameter configurations and architectural specifications utilized for the empirical evaluation of ReflecSched and the corresponding baselines.
Table H.1: Hyperparameter configurations for the baselines and the proposed ReflecSched.
Algorithm Parameter Value
IDDQN Training Episodes 2,000
Learning Rate / Discount Factor ($\gamma$) 0.01 / 0.9
Batch Size ($B$) / Buffer Size $O_{sum}$ / $10\times O_{sum}$
Exploration / PER ($\alpha,\beta$) $\epsilon$-greedy / 0.6, 0.4
Target Network Update Interval Every $O_{sum}$ steps
PPO-OC / DAN Learning Rate / Clipping Rate ($\epsilon$) $3\times 10^{-4}$ / 0.2
Value ($c_{v}$) / Entropy ($c_{e}$) Loss Coef. 0.5 / 0.01
GAE ($\lambda$) 0.95 (PPO), 0.98 (DAN)
PPO Update Epochs ($K$) 10 (PPO), 4 (DAN)
HMPSAC Training Episodes 1,000
Learning Rate / Discount Factor ($\gamma$) $3\times 10^{-4}$ / 0.99
Value ($c_{v}$) / Entropy ($c_{e}$) Loss Coef. 0.5 / 0.01
Max Decision Steps per Episode 10,000
ReflecSched Inference Temperature ($T_{inf}$) / Top-$p$ 0.2 / 0.8
Top-$k$ / Max Tokens 20 / 8,192
Reasoning Protocol Reflective (Non-thinking)
Rollout Expansion Factor ($N_{roll}$) 24
Maximum Search Depth ($L_{max}$) 6
Search Iterations ($N_{iter}$) 3
Search Sampling Temperature ($T_{search}$) 0.8
Dynamic Gantt Information / Static Data Enabled / Disabled